OpenStudy (yisselluna):

Find the quotient. n 12 ÷ n 4

OpenStudy (yisselluna):

help plzzzz!!!!

563blackghost (563blackghost):

Is you question.. \(\huge\bf{\frac{n^{12}}{n^{4}}}\) ?

OpenStudy (yisselluna):

n^12÷n^4

563blackghost (563blackghost):

\(\huge\bf{\frac{x^{m}}{x^{n}}=x^{m-n}}\)

OpenStudy (yisselluna):

i got n^8 but its wrong

563blackghost (563blackghost):

Hmmm it should be `n^8`. \(\huge\bf{\frac{n^{12}}{n^{4}}=n^{12-4}=n^{8}}\)

563blackghost (563blackghost):

@Directrix @3mar it should be `n^8` right?

OpenStudy (3mar):

\[\Huge\color{Fuchsia }\checkmark\]